OpenStudy (anonymous):

Find the value of the y-intercept of the line whose equation is 10x + 9y = 45.

OpenStudy (anonymous):

Simply get Y by itself on the left side of the equal sign.

OpenStudy (anonymous):

Slope = -2.222/2.000 = -1.111 x-intercept = 45/10 = 9/2 = 4.50000 y-intercept = 45/9 = 5